OpenStudy (anonymous):

This one I am stuck... Find the x-intercept of the equation of the line: -2y + 16x = -4.

OpenStudy (anonymous):

My mind is saying 4

OpenStudy (anonymous):

oops.. -4

OpenStudy (anonymous):

for x int put y=0

OpenStudy (anonymous):

okay well we have to put this in y=mx+b form -2y + 16x = -4. -2y=-16x-4 y=8x+2 now we set y=0 and solve for x 0=8x+2 -2=8x x=-.25 -.25 is the x-intercept

OpenStudy (anonymous):

no brent

OpenStudy (anonymous):

are u kidding me, i got a warning for my steps?? this is rediculous

OpenStudy (anonymous):

if u want x int put y=0 in eq x int is -1/4

OpenStudy (amistre64):

brents method is fine

OpenStudy (anonymous):

-1/4 is the same as -.25 last i checked...

OpenStudy (anonymous):

Thank you - figures it would be the fraction.... :O)

OpenStudy (anonymous):

how is it same

OpenStudy (amistre64):

\[\frac{1}{4}*\frac{25}{25}=\frac{25}{100}=.25\]